# Video Depth without Video Models

### v1

@@ Abstract
Figure 1: The RollingDepth model takes an unconstrained video and reconstructs a corresponding depth video. Unlike methods that rely on video diffusion models, it extends a single-image monodepth estimator such that it can process short snippets. To account for temporal context, snippets with varying frame rate are sampled from the video, processed, and reassembled through a global alignment algorithm to obtain long, temporally coherent depth videos. Depth is colour-coded from near Refer to caption far.
@@ 1 Introduction
Rather than design more refined video LDMs, which require huge resources for training, we take a step back and re-examine how far one can take video depth estimation with augmented single-image LDMs. We design a set of measures that, taken together, extend a per-image monodepth framework like Marigold [24] in a way that enables it to handle video input. Importantly, these measures greatly improve local and global consistency across time, while maintaining a constant memory footprint such that one can process long sequences. Specifically, we employ a “rolling” inference with a sliding window of a few frames (typically three, but other numbers are possible). Those snippets are sampled from the video with varying spacing, i.e., they can be immediately adjacent but also dilated along the timeline to cover long-range context. They are then fed into a multi-frame LDM fine-tuned from a single-frame model, with a modified cross-frame self-attention mechanism to enable information exchange. To reassemble the snippets, we propose a robust optimization-based global co-alignment, followed by averaging the aligned frames. Optionally, the resulting video can be degraded with moderate random noise and denoised again with the same per-snippet LDM to further refine spatial details.
@@ 2.1 Monocular Depth Estimation
Monocular depth estimation is a dense regression task. The pioneering work by Eigen et al. [12] showed that metric depth values can be recovered from single sensors. Successive advancements include including various parameterizations (ordinals, bins, planar guidance maps, piecewise planarity, CRFs, etc.[13, 30, 75, 47, 38, 2, 32, 44], switching CNN backbones to vision transformers [66, 33, 1, 3], considering camera intrinsics [74, 22, 19, 48], and patch-wise processing [35, 36, 5]. To handle “in-the-wild” settings, extensive internet photo collections are used for training [31, 73]. MiDaS [50] improves the generality by training on a mixture of multiple datasets. Depth Anything [68, 69] takes data scaling to the next level by relying on DINOv2 [45], a foundational model trained on 142M images in a self-supervised manner, and subsequently training with 62M pseudo-labels, 1M real depth annotations, and 0.5M synthetic ones. Recent trends leverage generative models, particularly diffusion models [57, 21], for depth estimation [80, 11, 53, 80]. Marigold [24] proposed to fine-tune Stable Diffusion [52], a generative text-to-image latent diffusion model (LDM) trained with LAION-5B [55], towards affine-invariant depth using 74k samples. This approach has been improved in many aspects including fewer steps [18, 16, 65, 20], finer details [77], and more modalities [14, 20].
@@ 2.2 Video Depth Estimation
Video depth estimation calls for dedicated mechanisms to ensure smoothness of adjacent frames, and correct handling of varying depth range. Existing approaches can be grouped into three main categories: test-time optimization, feed-forward prediction, and diffusion-based. Test-time optimization methods [42, 28, 78, 7] often rely on camera poses or optical flow and perform optimization for each new video during inference. While these methods can produce depth estimates that are temporally consistent, their dependence on camera poses and long processing time hamper their application to open-world video scenarios. Feed-forward prediction methods estimate depth sequences directly from input videos [34, 59, 76, 61, 72]. For example, DeepV2D [59] integrates camera motion estimation with depth prediction, MAMO [72] adopts memory attention mechanisms, and NVDS [62, 63] introduces a stabilization network as a post-processing module. However, the generalization of these methods to in-the-wild videos is often constrained by the limited diversity of training data and model capacity.
@@ 2.3 Image Diffusion Models for Video Tasks
Image diffusion models have been employed in various video inverse problems, such as video generation, inpainting, and super-resolution [29, 10]. A large amount of work [49, 70, 79] focusses on video editing, either by fine-tuning text-to-image diffusion models on video data [64, 39] or through training-free approaches using cross-frame attention and latent fusion [6, 26]. However, these works [64, 39, 6] predominantly address video-to-video translation tasks, where both the input and output reside in RGB space. In contrast, our approach leverages image diffusion priors to generate consistent depth videos, with the additional challenge to accommodate large variations of the depth range, as the near and far planes change – often suddenly – due to camera and object motion. Implementation tricks when using single-image models, like fixing the initial noise or blending consecutive latent representations, can somewhat mitigate the lack of knowledge w.r.t. temporal coherence, but do not solve it [24].
@@ 3.3 From Snippets to Video
Our multi-frame depth estimator operates on short snippets $n$ frames, where $n\ll N_{F}$. As snippets are processed independently, each has its own scale and shift – which are arbitrary in the case of affine-invariant methods [50, 24, 18] including Marigold, but will in practice not be aligned perfectly even when using a metric depth estimator [74, 48, 36]. To address that ambiguity, we construct overlapping snippets with different dilation rates. The frames shared between different snippets are subsequently used to align all depth predictions to a common scale and shift.
@@
Depth Co-alignment. At this stage we have generated $N_{T}$ depth snippets. Each of them has its own scale and shift parameters {$(s_{k},t_{k})$, $k\in 1\ldots T\}$, which are shared across its constituent frames. Our goal is to jointly compute the $N_{T}$ scale and shift values such that they optimally align all snippets into a consistent video. At a given frame $\mathbf{x}^{i}$, there are $N^{i}$ different individual depth maps $\{0pt^{i}_{j},j=1\ldots N^{i}\}$ originating from different snippets, where $N^{i}$ can vary from frame to frame. Let $k(i,j)$ be an indexing function that retrieves the snippet index $k$ for the $j$-th depthmap at frame $i$. To estimate the best alignment we minimize the $L_{1}$ loss over all individual depth predictions across depth predictions,
@@ 4.1 Implementation Details
Training Datasets. To finetune the snippet LDM, we use TartanAir [60], a synthetic video dataset with various (indoor and outdoor) scenes, styles, and camera motions. We visually inspect the scenes and select 18 scenes consisting of 369 sequences. Training snippets are randomly sampled from a sequence, with a minimum overlap ratio of 30%. To increase the diversity of scenes and avoid a significant sim-to-real gap we additionally also use Hypersim [51], a photorealistic single-image dataset containing 365 diverse scenes, treating images as 1-frame snippets.
@@ 4.2 Evaluation
Evaluation Datasets. We evaluate RollingDepth on four datasets that include both static and dynamic scenes with varying camera and scene motions: PointOdyssey [81] is a synthetic dataset with individually animated characters that move independently, designed for long-term tracking. We filter out overly simplistic toy scenes and retain 35 sequences. For each sequence, we follow the videodepth literature and exclude frames with camera zoom, then select the first 250 frames in each sequence. ScanNet v2 [8] is an indoor dataset of static scenes recorded with the Kinect RGB-D sensor. We use its test set of 100 sequences, downsample the frame rate by a factor 3, then keep the first 270 frames of each sequence. Bonn RGBD [46] is an RGB-D dataset of moving people in indoor spaces, captured with the OptiTrack Prime13 sensor. Following [23], we use frames 30-140 from five different dynamic scenes. DyDToF [58] is a photorealistic synthetic dataset featuring moving objects including people and animals. It has several videos per scene, we always take the first video and create two subsets of different lengths from it, by clipping frames 50-250, respectively frames 50-150.
Evaluation Protocol. We extend the affine-invariant depth evaluation protocol [50] to videos, i.e., depth predictions $\hat{0pt}$ are aligned to the ground truth with a scale and shift found with least squares fitting, where we fit one pair of transformation parameters per video, i.e., all frames in a video share a common scale and shift. We quantify the depth estimation accuracy with two standard metrics [50, 24, 62, 23]: the absolute mean relative error (AbsRel), defined as $\frac{1}{M}\sum_{j=1}^{M}{|\hat{0pt_{j}}-0pt_{j}|}/{0pt_{j}}$, where $M$ is the total number of pixels; and the $\delta 1$-accuracy, which measures the fraction of pixels for which $\max(\hat{0pt_{j}}/0pt_{j},0pt_{j}/\hat{0pt_{j}})<1.25$. Metrics are always given as percentages.
@@ 4.3 Comparison with Other Methods
Qualitative Comparison. To make our findings more tangible, we provide qualitative comparisons both on evaluation data and on in-the-wild examples. Fig. 4 confirms that RollingDepth consistently produces high-quality depth maps that preserve fine detail, both near the camera in the distance. DepthCrafter and ChronoDepth produce locally smooth videos with little frame-to-frame flicker, but have a tendency to distort the overall scene layout in a way that certain objects are segmented well but placed at incorrect (relative) depths. Single-image estimators are seemingly more accurate in that respect, but suffer from flickering and a lack of temporal coherence. We further illustrate these trends in Fig. 5, where we plot per-frame errors, as well as per-pixel errors for selected frames.
@@ A.1 Depth Co-Alignment
As discussed in Section 3.5 in the main paper, let $k(i,j)$ denote an indexing function that returns the snippet index $k$ corresponding to the $j$-th depthmap of $i$-th frame. To make the optimization more robust, we include an additional loss term in depth space while predicting inverse depth. We further scale the loss terms by their respective mean absolute value per frame to increase the numerical stability. Additionally, soft constraints on $s_{k},t_{k}$ are applied:
@@ A.2 Additional Training and Inference Details
During training, we apply gradient accumulation to increase the effective batch size, to 32. To better mix the samples with varying snippet lengths, every mini-batch is sampled randomly and can have different snippet lengths.
For the initial depth prediction, we apply the same random Gaussian noise to all frames. When applying refinement, the same noise is used to perturb the (encoded) co-aligned depth sequence. The denoising process then starts from timestep $T/2$.
@@ A.4 Baseline Methods
We evaluate baseline methods using their recommended default settings. For DepthCrafter [23], the inference is performed with 25 diffusion steps, using an overlap of 25 frames for videos longer than 110 frames. For ChronoDepth [56], inference comprises 10 denoising steps, with a window size of 10 (referred to as “num-frames” in the code) and a stride of 9 (referred to as “denoise-steps” in the code).
@@ Appendix B Additional Experiment Results
B.1 Failure cases of video models on PointOdyssey
We provide further examples from the PointOdyssey dataset where video-based methods struggle. Fig. S3 features scenes with large depth changes, such as hand gestures in front of the camera or objects entering the near field. These sudden changes require rapid alterations of the depth range, both before and after the event. Video models tend to produce incorrect overall scene layout in such cases, we hypothesize that they ”try too hard” to equalize the depth range throughout the scene.
B.2 Failure Cases of RollingDepth

### v2

@@ Abstract
Figure 1: The RollingDepth model takes an unconstrained video and reconstructs a corresponding depth video. Unlike methods that rely on video diffusion models, it extends a single-image monodepth estimator such that it can process short snippets. To account for temporal context, snippets with varying frame rates are sampled from the video, processed, and reassembled through a global alignment algorithm to obtain long, temporally coherent depth videos. Depth is colour-coded near Refer to caption far.
@@ 1 Introduction
Rather than design more refined video LDMs, which require huge resources for training, we take a step back and re-examine how far one can take video depth estimation with augmented single-image LDMs. We design a set of measures that, taken together, extend a per-image monodepth framework like Marigold [25] in a way that enables it to handle video input. Importantly, these measures greatly improve local and global consistency across time while maintaining a constant memory footprint such that one can process long sequences. Specifically, we employ a “rolling” inference with a sliding window of a few frames (typically three, but other numbers are possible). Those snippets are sampled from the video with varying spacing, i.e., they can be immediately adjacent but also dilated along the timeline to cover long-range context. They are then fed into a multi-frame LDM fine-tuned from a single-frame model, with a modified cross-frame self-attention mechanism to enable information exchange. To reassemble the snippets, we propose a robust optimization-based global co-alignment, followed by averaging the aligned frames. Optionally, the resulting video can be degraded with moderate random noise and denoised again with the same per-snippet LDM to further refine spatial details.
@@ 2.1 Monocular Depth Estimation
Monocular depth estimation is a dense regression task. The pioneering work by Eigen et al. [12] showed that metric depth values can be recovered from single sensors. Successive advancements include including various parameterizations (ordinals, bins, planar guidance maps, piecewise planarity, CRFs, etc.[13, 31, 79, 48, 39, 2, 33, 45, 85], switching CNN backbones to vision transformers [69, 34, 1, 3], considering camera intrinsics [78, 23, 20, 49, 50], and patch-wise processing [36, 37, 5]. To handle “in-the-wild" settings, extensive internet photo collections are used for training [32, 77]. MiDaS [52] improves the generality by training on a mixture of multiple datasets. Depth Anything [71, 72] takes data scaling to the next level by relying on DINOv2 [46], a foundational model trained on 142M images in a self-supervised manner, and subsequently training with 62M pseudo-labels, 1M real depth annotations, and 0.5M synthetic ones. Recent trends leverage generative models, particularly diffusion models [59, 22], for depth estimation [84, 11, 55, 84]. Marigold [25] proposed to fine-tune Stable Diffusion [54], a generative text-to-image latent diffusion model (LDM) trained with LAION-5B [57], towards affine-invariant depth using 74k samples. This approach has been improved in many aspects including fewer steps [18, 16, 67, 21], finer details [81], and more modalities [14, 21].
@@ 2.2 Video Depth Estimation
Video depth estimation calls for dedicated mechanisms to ensure smoothness of adjacent frames, and correct handling of varying depth range. Existing approaches can be grouped into three main categories: test-time optimization, feed-forward prediction, and diffusion-based. Test-time optimization methods [43, 29, 82, 7] often rely on camera poses or optical flow and perform optimization for each new video during inference. While these methods can produce depth estimates that are temporally consistent, their dependence on camera poses and long processing time hamper their application to open-world video scenarios. Feed-forward prediction methods estimate depth sequences directly from input videos [35, 61, 80, 63, 75, 76]. For example, DeepV2D [61] integrates camera motion estimation with depth prediction, MAMO [75] adopts memory attention mechanisms, and NVDS [64, 65] introduces a stabilization network as a post-processing module. However, the generalization of these methods to in-the-wild videos is often constrained by the limited diversity of training data and model capacity.
@@ 2.3 Image Diffusion Models for Video Tasks
Image diffusion models have been employed in various video inverse problems, such as video generation, inpainting, and super-resolution [30, 10, 86]. A large amount of work [51, 73, 83] focusses on video editing, either by fine-tuning text-to-image diffusion models on video data [66, 40] or through training-free approaches using cross-frame attention and latent fusion [6, 27]. However, these works [66, 40, 6] predominantly address video-to-video translation tasks, where both the input and output reside in RGB space. In contrast, our approach leverages image diffusion priors to generate consistent depth videos, with the additional challenge to accommodate large variations of the depth range, as the near and far planes change – often suddenly – due to camera and object motion. Implementation tricks when using single-image models, like fixing the initial noise or blending consecutive latent representations, can somewhat mitigate the lack of knowledge w.r.t. temporal coherence, but do not solve it [25].
@@ 3.3 From Snippets to Video
Our multi-frame depth estimator operates on short snippets of $n$ frames, where $n\ll N_{F}$. As these snippets are processed independently, each has its own scale and shift – which are arbitrary in the case of affine-invariant methods [52, 25, 18] including Marigold, but will in practice not be perfectly aligned even when using a metric depth estimator [78, 49, 37]. To resolve that ambiguity, we construct overlapping snippets with different temporal dilation rates. The frames shared between different snippets are subsequently used to align all depth predictions to one common scale and shift.
@@
Depth Co-alignment. At this stage we have generated $N_{T}$ depth snippets. Each of them has its own scale and shift parameters {$(s_{k},t_{k})$, $k\in 1\ldots T\}$, which are shared across its constituent frames. Our goal is to jointly compute $N_{T}$ scale and shift values such that they optimally align all snippets into a consistent video. At a given frame $\mathbf{x}^{i}$, there are $N^{i}$ different individual depth maps $\{0pt^{i}_{j},j=1\ldots N^{i}\}$ originating from different snippets, where $N^{i}$ can vary from frame to frame. Let $k(i,j)$ be an indexing function that retrieves the snippet index $k$ for the $j$-th depthmap at frame $i$. To estimate the best alignment, we minimize the $L_{1}$ loss over all individual depth predictions,
@@ 4.1 Implementation Details
Training Datasets. To finetune the snippet LDM, we use TartanAir [62], a synthetic video dataset with various (indoor and outdoor) scenes, styles, and camera motions. We visually inspect the scenes and select 18 scenes consisting of 369 sequences. Training snippets are randomly sampled from a sequence, with a minimum overlap ratio of 30%. To increase the diversity of scenes and avoid a significant sim-to-real gap we additionally use Hypersim [53], a photorealistic single-image dataset containing 365 diverse scenes, treating images as 1-frame snippets.
@@ 4.2 Evaluation
Evaluation Datasets. We evaluate RollingDepth on four datasets that include both static and dynamic scenes with varying camera and scene motions: PointOdyssey [87] is a synthetic dataset with individually animated characters that move independently, designed for long-term tracking. We filter out overly simplistic toy scenes and retain 35 sequences. For each sequence, we follow the videodepth literature and exclude frames with camera zoom, then select the first 250 frames in each sequence. ScanNet v2 [8] is an indoor dataset of static scenes recorded with the Kinect RGB-D sensor. We use its test set of 100 sequences, taking the first 270 frames of each sequence and downsample the frame rate by a factor 3, Bonn RGBD [47] is an RGB-D dataset of moving people in indoor spaces. Following [24], we use frames 30-140 from five different dynamic scenes. DyDToF [60] is a photorealistic synthetic dataset featuring moving objects including people and animals. It has several videos per scene, we always take the first video and create two subsets of different lengths from it, by clipping frames 50-250, respectively frames 50-150.
Evaluation Protocol. We extend the affine-invariant depth evaluation protocol [52] to videos, i.e., depth predictions $\hat{0pt}$ are aligned to the ground truth with a scale and shift found with least squares fitting, where we fit one pair of transformation parameters per video, i.e., all frames in a video share a common scale and shift. We quantify the depth estimation accuracy with two standard metrics [52, 25, 64, 24]: the absolute mean relative error (AbsRel), defined as $\frac{1}{M}\sum_{j=1}^{M}{|\hat{0pt_{j}}-0pt_{j}|}/{0pt_{j}}$, where $M$ is the total number of pixels; and the $\delta 1$-accuracy, which measures the fraction of pixels for which $\max(\hat{0pt_{j}}/0pt_{j},0pt_{j}/\hat{0pt_{j}})<1.25$. Metrics are always given as percentages. We provide additional temporal smoothness evaluation in the supplementary material.
@@ 4.3 Comparison with Other Methods
Qualitative Comparison. To make our findings more tangible, we provide qualitative comparisons both on evaluation data and on in-the-wild examples. Figure. 4 confirms that RollingDepth consistently produces high-quality depth maps that preserve fine detail, both near the camera in the distance. DepthCrafter and ChronoDepth produce locally smooth videos with little frame-to-frame flicker, but have a tendency to distort the overall scene layout in a way that certain objects are segmented well but placed at incorrect (relative) depths. Single-image estimators are seemingly more accurate in that respect, but suffer from flickering and a lack of temporal coherence. We further illustrate these trends in Fig. 5, where we plot per-frame errors, as well as per-pixel errors for selected frames.
@@ A.1 Depth Co-Alignment
As discussed in Sec. 3.3, let $k(i,j)$ denote an indexing function that returns the snippet index $k$ corresponding to the $j$-th depthmap of $i$-th frame. To make the optimization more robust, we include an additional loss term in depth space while predicting inverse depth. We further scale the loss terms by their respective mean absolute value per frame to increase the numerical stability. Additionally, soft constraints on $s_{k},t_{k}$ are applied:
@@ A.2 Additional Training and Inference Details
During training, we follow Marigold to use MSE loss on the latents. We apply gradient accumulation to increase the effective batch size, to 32. To better mix the samples with varying snippet lengths, every mini-batch is sampled randomly and can have different snippet lengths. For the initial depth prediction, we apply the same random Gaussian noise to all frames. When applying refinement, the same noise is used to perturb the (encoded) co-aligned depth sequence. The denoising process then starts from timestep $T/2$.
@@ A.4 Baseline Methods
We evaluate baseline methods using their recommended default settings. For DepthCrafter [24], the inference is performed with 25 diffusion steps, using an overlap of 25 frames for videos longer than 110 frames. For ChronoDepth [58], inference comprises 10 denoising steps, with a window size of 10 (referred to as “num-frames" in the code) and a stride of 9 (referred to as “denoise-steps" in the code).
@@ Appendix B Additional Experiment Results
B.1 Temporal smoothness evaluation
We further quantitatively evaluate the temporal smoothness using optical-flow-based warping loss (OPW) [63] on PointOdyssey and ScanNet datasets and report the results in Tab. S1. The optical flow is estimated using GMFlow [68].
Table S1: Temporal smoothness (OPW↓) comparison. All values are ×$\times 10^{3}$, lower is better. denotes catastrophic failures on some sequences. Numbers in brackets are evaluated on subsets that exclude those cases.
PointOdyssey ScanNet
Marigold 3.52 (4.00) 0.48
DepthAnything 3.92 (4.21) 0.32
NVDS 3.50 (2.97) 0.29
ChronoDepth 8.98 (2.99) 0.29
DepthCrafter 7.75 (1.30) 0.25
RollingDepth (ours) 1.42 (1.63) 0.20
We notice that ChronoDepth and DepthCrafter have catastrophic failure in some cases of PointOdyssey (cf. Sec. B.4), leading to large errors, as denoted by . We manually exclude these failure cases. The re-calculated average OPW is reported in the brackets. Overall, RollingDepth shows good smoothness, on par with DepthCrafter, while being more robust than DepthCrafter and ChronoDepth against occasional failures.
We point out that OPW only evaluates the “smoothness” between adjacent frames while ignoring the long-term smoothness and geometric consistency. As shown in Tab. S2, with larger dilation rates, the geometric accuracy shows a clear improving trend, while the trend of OPW is unclear. We hypothesize that with a larger dilation rate, geometric accuracy is improved at a cost of minor local smoothness decrease when merging the aligned snippets.
Table S2: Extended table of dilation rate ablation study (Tab. 2). Values are ×$\times 10^{3}$.
PointOdyssey ScanNet
Dilation rates Abs Rel↓ $\delta 1$ OPW ↓ Abs Rel↓ $\delta 1$ OPW ↓
{1} 16.7 75.5 1.22 12.8 83.2 0.24
{1, 25} 10.2 89.5 2.06 10.6 88.8 0.29
{1, 10, 25} 10.2 89.6 1.98 9.9 90.1 0.29
B.2 Evaluation on DDAD dataset
We further evaluate the model performance on the DDAD [19] dataset, which is a driving-scene dataset with sparse depth annotation. We use the 100-frame sequences on the test set.
As shown in Tab. S3, RollingDepth outperforms other methods in terms of accuracy and smoothness.
Table S3: Evaluation on DDAD dataset.
Abs Rel↓ $\delta 1$ OPW ↓
×$\times 10^{-2}$$\times 10^{-2}$$\times 10^{-3}$
NVDS 30.8 57.2 0.39
ChronoDepth 34.2 46.9 0.21
DepthCrafter 19.3 74.8 0.28
RollingDepth (ours) 12.8 83.2 0.19
B.3 Inference efficiency
We report the inference efficiency comparison in Tab S4. The benchmarking is done on the same machine with a single RTX3090 GPU. For each method, we run 10 repeated inferences after a warm-up iteration, with the model loaded on GPU, and calculated the mean run time and peak memory footage of each iteration.
Table S4: Inference speed and peak GPU memory usage comparison on a 768$\times$432 video of 250 frames. By increasing the batch size of processing, $\text{RollingDepth}^{\dagger}$ can trade memory for speed.
Time (s) Peak GPU Memory (GB)
NVDS 284 17.6
ChronoDepth 121 15.0
DepthCrafter 284 13.6
RollingDepth (ours) 105 16.2
$\text{RollingDepth}^{\dagger}$ (ours) 181 40.1
B.4 Failure cases of video models on PointOdyssey
We provide further examples from the PointOdyssey dataset where video-based methods struggle. Figure S2 features scenes with large depth changes, such as hand gestures in front of the camera or objects entering the near field. These sudden changes require rapid alterations of the depth range, both before and after the event. Video models tend to produce incorrect overall scene layout in such cases, we hypothesize that they "try too hard" to equalize the depth range throughout the scene.
B.5 Failure Cases of RollingDepth